Latest Headlines:

Sol Summary – November 18th

November 18, 2016

Sol Summary Report (SSR):

Sol# 56

Person filling out report: Annalea Beattie

Summary Title: Chilly Friday.

Mission Status: On track

Sol Activity Summary:

At the start of our day, Plant Number Seven pH is down. O no.

Claude-Michel advises it needs the seeds pushing in one quarter of an inch and ten drops of pH.  We hope this works.

Early, Anu and Yusuke went out on EVA to sample lichens, I asked, but there is nothing more to say except a successful sampling EVA and an early finish due to helmet fogging. Anastasiya  and Claude-Michel partially assembled the new ATV in the morning and Yusuke helped finish it this afternoon. Jon and Annalea took ATVs to a canyon we all call Green Mars for a very chilly EVA to sample hypoliths. In the afternoon Anu and Anastasiya worked in the laboratory on hypolith processing. I cooked a loaf of bread which looks and tastes like sandstone. We all ate chocolate and chips with some masala tea mid arvo and after dinner tonight THE EXPANSE!

Reports Submitted to CapCom:

  1. Sol Summary- Annalea
  2. EVA time lapse- Yusuke
  3. Journalist report – Anastasiya
  4. Alex- Mars 160 rocks.
  5. Pictures -YES
  6. Photo of the Day -YES

MDRS lessons:

Helmets fog more in cool weather.

Plans for tomorrow:  Follow-up EVAs, similar but not the same.

Crew Physical Status: Very well, how are you?

Weather: Minus seven overnight, now hot.

Anomalies: Helmet fog is unusual but contagious in cold weather.